OpenStudy (anonymous):

WILL FAN AND MEDAL Check my answer. What is the slope of the line with equation y = –3? A. –3 B. 0 C. 3 D. The slope is undefined. Is it D?

OpenStudy (hoblos):

y=mx+b where m is the slope here, y=-3 so y= 0x - 3 so what would be the slope ?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

I struggles with this in class today

OpenStudy (anonymous):

I didn't understand the way she taught it but I sort of made out what I could understand

OpenStudy (anonymous):

struggled*

OpenStudy (hoblos):

compare (y=mx+b) and (y=0x-3) which number corresponds to 'm' in the second equation

OpenStudy (hoblos):

@superwholockedgirl you still not getting it ?

OpenStudy (anonymous):

no so is it a b c or d

OpenStudy (hoblos):

in y=0x-3 what is the number multiplying x ?

OpenStudy (anonymous):

idk

OpenStudy (jdoe0001):

 \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ -3}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ -3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\) so.... what would that give you as the slope ?

OpenStudy (anonymous):

this is all they give me.

OpenStudy (anonymous):

What is the slope of the line with equation y = –3? A. –3 B. 0 C. 3 D. The slope is undefined.

OpenStudy (jdoe0001):

so, try finding the slopes with those 2 points on the line y = -3 what would you get?

OpenStudy (anonymous):

It's D. I went to Yahoo answers

OpenStudy (hoblos):

it's not D

OpenStudy (anonymous):

My mom said it was D, too

OpenStudy (anonymous):

I have like 3+ people saying it's D.

OpenStudy (hoblos):

well we are trying to show you how to solve it by yourself and get your own correct answer.. to ways are shown, try using one of them. and it's absolutely not D